From GlueXWiki
Jump to: navigation, search

November 7, 2013 FDC meeting


  1. Installation status[1]
  2. Engineering (Bill)
  3. Electronics (Chris, Nick)
  4. Other


Participants: Eugene, Fernando, Dave, Nick, Simon, Vlad, Beni, Bill(came later), and Lubomir.


- FDC has been moved on the platform and inserted in the magnet last Friday. Inside the magnet a survey of the downstream and upstream side has been done. Also FDC has been moved inside the magnet, and re-surveyed. Repeatability was very good (except for z-direction translation). Once the reflectors are installed the FDC will be moved again to the floor to survey the reflectors w.r.t. the already surveyed points. The work on the mesh has not started yet.

- Bill received one reflector (made out of fused silica) and glued it to the holder. It has been tested then by the survey group: they found that it can be used at up to 6 m distance from the Fero unit with a resolution of 0.1 mm. The resolution was estimated from the repeatability of the measurements, keeping the positions the same. The new 16 reflectors (regular glass) have arrived and Bill will glue them to the holders. He will give one to the survey group for tests again, since the material is different. At the beginning of next week he will glue the holders to the gusset rings of the packages; there's a jig to fix their positions.

- The chilled water system must be ready now. According to Bill it will take just several hours to connect the chiller to the FDC and the chilled water, then we can test the FDC for leakage.


- Thickness of the packages: the thickness of a cell used so far in the simulation and data analyses was 2.1254 cm. By comparing the angles reconstructed from the Indiana chambers to those from the FDC, Lubomir estimated the thickness should be 2.15 +/- 0.1 cm (for the third package). This created some confusion that the nominal value may differ from the measured one due to the thickness of the glue on the cathodes that is not well controllable. However the problem was solved when Bill looked at the recent I-deas model and found the value there's 2.145 cm (after the meeting, xml file attached). He calculated also from the model the thickness of the full package: 14.42 cm and compared it to the results of the gusset flatness survey (attached) which gives 14.43 cmm for the first package: very good agreement! Bill is going to look at the other packages too.


- The LV system will be installed at the north side; right now only the patch panels are there but Nick will install the power supplies as well. The HV system is already installed. After some discussions, the best will be to first install all the LV cables, route all of them to the top north cable tray and connect them permanently to the panels there. The cables that come from the south side should be lifted above the mesh and then directed to the north: we need open views to the reflectors.

- At the beginning one VXS crate will be installed (even temporary) at the north side to be used in the mini-DAQ to test the cable connections during the first phase of the cabling. It will have at least one F1TDC and one fADC125. For the crate and the LV power supplies we may need 208V and therefore we will move the transformer there. We expect the racks to be powered in several weeks from now.

- Nick now has all the LV and signal cables, except for the spare signal cables for package 1.